# Arnold-Thom conjecture for the arrival time of surfaces

### v1

@@ 4.3. Uniqueness of limit tangents
We will finish the proof of the Arnold-Thom conjecture for non-degenerate singularities in this section.
Theorem 4.23 (Arnold-Thom for non-degenerate singularities).
@@
The uniqueness of limit points is addressed in Theorem 3.1. For the uniqueness of limit tangents, by Theorem 4.4, at a non-degenerate cylindrical singularity, all the limit flows are the same type of cylinders with parallel axes. Thus, the same arguments in the proof of the second case of Theorem 3.11 apply and the conclusion follows. ∎
@@ 5.1. Outer flows
Note that, a priori, we know that $\textbf{H}_{M_{t}}$ is defined almost everywhere. This allows us to have the definition above, but we do not have any existence and uniqueness theorems as in the usual ordinary differential equation theory. In fact, a flow line can have strange behavior, especially at those points where the mean curvature of the flow is not defined. However, the mean convex neighborhood theorem [CHHW, Theorem 9.1] implies that near an inward neck singularity, the flow is mean convex. Therefore, close to the neck singularity, $\textbf{H}_{M_{t}}\left(\gamma(t)\right)\neq 0$ for $t<\infty,$ so we can revert to the classical setting when approaching a neck singularity. The rest of the proofs in Section 3 are valid, and we get the Łojasiewicz theorem and Arnold-Thom conjecture in the case of outer flows.
@@
Let $M_{t}$ be the outer flow starting from a closed smooth hypersurface $M$ in $\mathbb{R}^{n+1}$ and let $(0,0)$ be an inward neck singularity of it. If $\gamma$ is a gradient flow line of $M_{t}$ with a limit point, then the limit point is unique and the unit tangents of $\gamma$ converge.
In fact, with the notion of flow lines defined above, the Łojasiewicz theorem and the Arnold-Thom conjecture also hold for a non-degenerate cylindrical singularity in a general mean curvature flow even without the mean convexity condition. See Remark 4.12.
We remark that gradient flow lines in the context of Theorem 5.3 may not be globally $C^{1}.$ However, the theorem only cares about the behaviors of flow lines near inward neck singularities. In a small neighborhood of an inward neck singularity, the flow is a mean convex level set flow so a gradient flow line can be interpreted in the classical sense. This again manifests the principle proposed by Colding-Minicozzi [CM19] that solutions with low regularity can behave like analytic functions.
@@ 5.2. Flows of two-spheres and generic hypersurfaces
We mention three situations in which the general framework developed in this section applies directly without the issue of fattening. The first case is when a flow starts from a two-sphere, and the second and the third cases are when a flow starts from a generic closed hypersurface.
With the works of [W95, Br16, BK23], level set flows starting from a two-sphere are understood well. In fact, this was already discussed in [CHH, Proof of Theorem 1.12] assuming the multiplicity one conjecture, which was recently proven by Bamler-Kleiner [BK23] (cf. [BK23, Corollary 1.3]). In particular, the level set flow starting from a two-sphere does not fatten. Thus, starting from a smooth two-sphere, its inner flow, its outer flow, and its level set flow will all be the same. Theorem 5.3 then implies the following uniqueness results (Łojasiewicz theorem and Arnold-Thom conjecture).
@@
The second situation is related to the study of generic mean curvature flows of surfaces. The program was initiated by Colding-Minicozzi [CM12], where they prove that generalized cylinders are the only linearly stable singularity models. The possibilities of getting generic singularities by perturbing initial data were confirmed in [CCMS, CCS, SX1, SX2], with some potential scenario of high multiplicities. Again, with the recent work of Bamler-Kleiner [BK23], combined with the work of Chodosh–Choi-Mantoulidis–Schulze [CCMS] and Chodosh–Choi–Schulze [CCS], it is now clear that mean curvature flows starting from generic closed embedded surfaces only develop cylindrical or spherical singularities. Thus, Theorem 5.3 again implies the following uniqueness results.
@@
We can also deal with generic low-entropy mean curvature flows in $\mathbb{R}^{n+1}$ with $n+1\in\left\{4,5,6\right\}$. By the work of Chodosh–Choi-Mantoulidis–Schulze [CCMS2] and Chodosh-Mantoulidis–Schulze [CMS], if the entropy of the initial hypersurface is low, then the generic mean curvature flow only encounters cylindrical singularities. In particular, Theorem 5.3 implies the following result when the entropy is not greater than that of a bubble sheet $S^{n-2}\times\mathbb{R}^{2}.$
@@ Appendix A Convergence of flow line
Thus, we prove that $T_{\varepsilon}=T,$ and hence the convergence of $\gamma^{i}$ on $[-T,0]$ follows from the locally smooth convergence of the flows. ∎

### v2

@@ 4.3. Uniqueness of limit tangents
We will finish the proof of the Arnold–Thom conjecture for non-degenerate singularities.
Theorem 4.22 (Arnold–Thom for non-degenerate singularities).
@@
The uniqueness of limit points is addressed in Theorem 4.13. For the uniqueness of limit tangents, by Theorem 4.4, at a non-degenerate cylindrical singularity, all the limit flows are the same type of cylinders with parallel axes. Thus, applying Lemma 3.10 finishes the proof. ∎
@@ 5.1. Outer flows
Note that, a priori, we know that $\textbf{H}_{M_{t}}$ is defined almost everywhere. This allows us to have the definition above, but we do not have any existence and uniqueness theorems as in the usual ordinary differential equation theory. In fact, a flow line can have strange behavior, especially at those points where the mean curvature of the flow is not defined. However, the mean convex neighborhood theorem [CHHW, Theorem 9.1] implies that near an inward neck singularity, the flow is mean convex. Therefore, close to the neck singularity, $\textbf{H}_{M_{t}}\left(\gamma(t)\right)\neq 0$ for $t<\infty,$ so we can revert to the classical setting when approaching a neck singularity. The rest of the proofs in Section 3 are valid, and we get the Łojasiewicz theorem and Arnold–Thom conjecture in the case of outer flows.
@@
Let $M_{t}$ be the outer flow starting from a closed smooth hypersurface $M$ in $\mathbb{R}^{n+1}$ and let $(0,0)$ be an inward neck singularity of $M_{t}.$ If $\gamma$ is a gradient flow line of $M_{t}$ with a limit point, then the limit point is unique and the unit tangents of $\gamma$ converge.
In fact, with Definition 5.2, the Łojasiewicz theorem and the Arnold–Thom conjecture also hold for a non-degenerate cylindrical singularity in a general mean curvature flow even without the mean convexity condition. See Remark 4.12.
We remark that gradient flow lines in the context of Theorem 5.3 may not be globally $C^{1}.$ However, the theorem only cares about the behaviors of flow lines near inward neck singularities. In a small neighborhood of an inward neck singularity, the flow is a mean convex level set flow so a gradient flow line can be interpreted in the classical sense. This again manifests the principle proposed by Colding–Minicozzi [CM19] that solutions with low regularity can behave like analytic functions.
@@ 5.2. Flows of two-spheres and generic hypersurfaces
We mention three situations in which the general framework developed in the previous section applies directly without the issue of fattening. All flows are assumed to be embedded. The first case is when a flow starts from a two-sphere, and the second and the third cases are when a flow starts from a generic closed hypersurface.
With the works of [W95, Br16, HW, CHH, BK23], level set flows starting from a two-sphere are understood well. In fact, this was already discussed in [CHH, Proof of Theorem 1.12] assuming the multiplicity one conjecture, which was recently proven by Bamler–Kleiner [BK23] (cf. [BK23, Corollary 1.3]). In particular, the level set flow starting from a two-sphere does not fatten. Thus, starting from a smooth two-sphere, its inner flow, its outer flow, and its level set flow will all be the same. Theorem 5.3 then implies the following uniqueness results (Łojasiewicz theorem and Arnold–Thom conjecture).
@@
The second situation is related to the study of generic mean curvature flows of surfaces. The program was initiated by Colding–Minicozzi [CM12], where they proved that generalized cylinders are the only linearly stable singularity models. The possibilities of getting generic singularities by perturbing initial data were confirmed in [CCMS, CCS, SX1, SX2], with some potential scenario of high multiplicities. Again, with the recent work of Bamler–Kleiner [BK23], combined with the work of Chodosh–Choi–Mantoulidis–Schulze [CCMS] and Chodosh–Choi–Schulze [CCS], it is now clear that mean curvature flows starting from generic closed embedded surfaces only develop neck or spherical singularities. Thus, Theorem 5.3 again implies the following uniqueness results.
@@
We can also deal with generic low-entropy mean curvature flows in $\mathbb{R}^{n+1}$ with $n+1\in\left\{4,5,6\right\}$. By the work of Chodosh–Choi–Mantoulidis–Schulze [CCMS2] and Chodosh–Mantoulidis–Schulze [CMS], if the entropy of the initial hypersurface is low, then the generic mean curvature flow only encounters cylindrical singularities. In particular, Theorem 5.3 implies the following result when the entropy is not greater than that of a bubble sheet $S^{n-2}\times\mathbb{R}^{2}.$
@@ Appendix A Convergence of flow line
Thus, we prove that $T_{\varepsilon}=T,$ and hence the convergence of $\gamma^{i}$ on $[-T,0]$ follows from the locally smooth convergence of the flows. This proves the convergence on any $[-T,0]\subseteq(-\overline{T},0].$
Appendix B Proof of Proposition 2.8
In this appendix, we give a streamlined proof of Proposition 2.8.
Proof of Proposition 2.8.
Without loss of generality, we may assume that $p=0$ is the origin and the unique tangent flow at $(0,0)$ is the neck $\Gamma=\big{\{}(x_{1},...,x_{n+1})\ |\ x_{1}^{2}+\cdots+x_{n}^{2}=2(n-1)\big{\}}$ with axis =$\ell=\big{\{}(x_{1},...,x_{n+1})\ |\ x_{1}=x_{2}=\cdots=x_{n}=0\big{\}}.$
By the definition of tangent flows, for each $\varepsilon,$ there is a constant $\rho=\rho(\varepsilon)>0$ with $\rho(\varepsilon)\rightarrow 0$ as $\varepsilon\rightarrow 0$ such that for all $0<r\leq\rho$ we have that
$\frac{1}{\sqrt{-t}}M_{t}^{(0,0,r^{-1})}$ is $\frac{\varepsilon}{4}$ close to $\Gamma$ when $t\in[-2,-\frac{1}{2}].$
We define $Z^{\rho(\varepsilon)}(\bar{x},\bar{t})$ to be
$Z^{\rho(\varepsilon)}(\bar{x},\bar{t}):=\inf\left\{z:\begin{aligned} &\frac{1}{\sqrt{-t}}M_{t}^{(\bar{x},\bar{t},r^{-1})}\text{ is }\varepsilon\text{-close to some neck in }\mathcal{C}_{1}^{0}\\ &\text{ for each }r\in(z,\rho(\varepsilon))\text{ and }t\in\left[-2,-1/2\right]\end{aligned}\right\}.$
We also define $Z_{i}=Z^{\rho(\varepsilon)}\left(\gamma(t_{i}),t_{i}\right)$. Since $(\gamma(t_{i}),t_{i})$ is regular point, it follows that $Z_{i}>0$. Now we rescale the flow by $Z_{i}^{-1}$ around $(\gamma(t_{i}),t_{i})$. Namely, we consider
$M^{i}_{t}:=M_{t}^{(\gamma(t_{i}),t_{i},Z_{i}^{-1})}.$
Then, for each $0<r\leq\rho(\varepsilon)$, there exists $i_{0}=i_{0}(r)$ such that for all $i\geq i_{0}$ we have that
$\frac{1}{\sqrt{-t}}M_{t}^{\left(\gamma(t_{i}),t_{i},s^{-1}\right)}\text{ is }\frac{\varepsilon}{2}\text{-close to }\Gamma\text{ for all }s\in[r,\rho(\varepsilon)]\text{ when }t\in\left[-2,-\frac{1}{2}\right].$
By definition, this implies that $\limsup\limits_{i\rightarrow\infty}Z_{i}\leq r$. Since $r$ is arbitrary, we have $\lim\limits_{i\to\infty}Z_{i}=0$ and $\lim\limits_{i\rightarrow\infty}\frac{\rho(\varepsilon)}{Z_{i}}=\infty.$ By the convergence theorem of White (c.f [W00, W03, HK1]), $M_{t}^{i}$ converges smoothly to a noncollapsed ancient solution $M_{t}^{\infty}$ such that $0\in M_{0}^{\infty}.$ By the classification result of [BC19, BC21], the definition of $Z$, and Lemma 2.6, we know that $M_{t}^{\infty}$ must be a shrinking neck or a bowl soliton with axis $\ell^{\prime}$. Moreover, $M^{\infty}_{t}$ is not $\frac{\varepsilon}{2}$-close to any neck in $\mathcal{C}_{1}^{0}$ for some $t\in\left[-2,-\frac{1}{2}\right]$.
We first claim that $\ell^{\prime}$ is at most $C\varepsilon$ away from $\ell$. (Namely, the angle between $\ell^{\prime}$ and $\ell$ is at most $C\varepsilon.$) In fact, by the definition of $Z$, we have $\overline{M}_{\tau}^{i}=e^{\frac{\tau}{2}}M_{-e^{-\tau}}^{i}$ is $\varepsilon$-close to $\Gamma$ for all $\tau\in\left[-\ln 2-\ln\frac{\rho(\varepsilon)}{Z_{i}},\ln 2\right]$. By Colding–Minicozzi’s Łojasiewicz inequality [CM15], the axis of the cylinder has total rotation at most $C\varepsilon$ during $\tau\in\left[-\ln\frac{\rho(\varepsilon)}{Z_{i}},0\right]$ (c.f. [GH, Proposition 4.1] and [CM15, Lemma A.48]). By definition of $M^{i}$ and $\rho(\varepsilon)$, $\overline{M}_{\tau}^{i}$ is $\varepsilon$-close to $\Gamma$ when $\tau=-\ln\frac{\rho(\varepsilon)}{Z_{i}}$ as long as $i$ is large enough. Therefore, $\overline{M}_{\tau}^{i}$ is $C\varepsilon$-close to $\Gamma$ for all $\tau\in\left[-\ln\frac{\rho(\varepsilon)}{Z_{i}},0\right]$. Passing to the limit proves the claim.
We next claim that $C^{-1}<H(\gamma(t_{i}),t_{i})Z_{i}\leq C$. If this is not the case, then passing to the limit, we have $H(0,0)=0$ or $H^{-1}(0,0)=0$ in $M^{\infty}_{t}$. The first option is not consistent with the only two possibilities of $M^{\infty}_{t}$. For the second case, $M^{\infty}_{t}$ must be a shrinking neck with extinction time at 0 and with axis passing through the origin 00. However, this contradicts the fact that $M^{\infty}_{t}$ is not $\frac{\varepsilon}{2}$-close to any neck in $\mathcal{C}_{1}^{0}$ for some $t\in\left[-2,-\frac{1}{2}\right]$.
As a consequence, any subsequential limit $\widetilde{M}^{\infty}_{t}$ of $M_{t}^{(\gamma(t_{i}),t_{i},H(\gamma(t_{i}),t_{i})^{-1})}$ must be a parabolic rescaling of $M_{t}^{\infty}$, i.e., $\widetilde{M}^{\infty}_{t}$ = $\lambda{M}^{\infty}_{\lambda^{-2}t}$ for some $\lambda\in[C^{-1},C]$. Hence, $\widetilde{M}^{\infty}_{t}$ is still a shrinking neck or bowl soliton with axis at most $C\varepsilon$ away from $\ell$.
Since $\varepsilon$ is arbitrary, we conclude that (after passing to a further subsequence) $\widetilde{M}^{\infty}_{t}$ is a shrinking neck or bowl soliton with its axis exactly $\ell$. This finishes the proof. ∎